OpenStudy (anonymous):

Write quadratic equation in standard form with given solution set: {1/6, -2/3}

OpenStudy (anonymous):

since 1/6 is a zero on factor is \[x-\frac{1}{6}\] another is \[x+\frac{2}{3}\] so you could write \[(x-\frac{1}{6})(x+\frac{2}{3})\]

OpenStudy (anonymous):

but instead you would probably write \[(6x-1)(3x+2)\] standard form means multiply this out to get \[18x^2+9x-2\] if my algebra is correct.

OpenStudy (anonymous):

And then what? Use the quadratic formula?

OpenStudy (anonymous):

no. you already have the zeros. you were given that the zeros are \[\frac{1}{6}\] and \[-\frac{2}{3}\] so you are not being asked to find the zeros. this question is asked in reverse. not "here is the equation, find the zeros" but rather "here are the zeros, find the equation"

OpenStudy (anonymous):

So the equation would be my final answer then, right?

OpenStudy (anonymous):

if you are given that the zeros are a and b then the equation must have been \[(x-a)(x-b)=0\]

OpenStudy (anonymous):

yes the equation is what you were asked for. if i multiplied out correctly, your answer should be \[18x^2+9x-2=0\]

OpenStudy (anonymous):

Awesome, thank you so much.

OpenStudy (anonymous):

welcome, but check my algebra, i am tired.

OpenStudy (anonymous):

It is correct.